OpenStudy (anonymous):

factor completely: 2x^2y-5xy-12y

OpenStudy (nowhereman):

wolframalpha is your friend ^^: \[(x-4)(2x+3)y\]

OpenStudy (anonymous):

2x^2y-5xy-12y = y(2x^2 -5x-12) = y(2x+3)(x-4)